# A Survey on Federated Analytics: Taxonomy, Enabling Techniques, Applications and Open Issues

### v1

@@ V-A Privatization techniques
Compared to LDP, where the client data are considered data samples and directly protected, CDP cannot be directly applied to many FA scenarios because it considers enforcing protection on the data already aggregated. In the FA literature, there are mainly two settings to enforce CDP. In the first setting, the raw data of each client is considered as a database consisting of multiple data samples. The clients enforce CDP on the database, so that information about each individual sample is protected. FA research considering this setting includes [86, 49]. In the second setting, similar to the setting of LDP, the data within a single client is considered as a data sample in $\mathcal{D}$. The CDP scheme guarantees that an adversary cannot infer the sensitive information of a client by analyzing the final FA result. However, such privacy attacks can succeed when the adversary has access to the uploads of the clients (e.g., when the server acts as the adversary). Therefore, the second setting usually assumes a trusted aggregator (server), and sacrifices the formal privacy guarantee. The second setting has been considered in [87, 88, 18, 43].
@@
These exist diverse techniques to realize CDP. Similar to those used for LDP, the Gaussian mechanism can still be applied for CDP to protect continuous scaler output, which can satisfy $(\epsilon,\delta)$-CDP [86, 49, 87, 88]. [^2] Laplace mechanism also has theoretically applicable for CDP, but is not selected for researchers due to the high variance of Laplace noise. Compared to the LDP case, where Gaussian noise is added to each single data sample, applying the Gaussian mechanism for CDP only requires adding noise into the aggregated result. Therefore, the overall noise on the final analytics results is significantly reduced. The exponential mechanism tackles the analytics tasks where a single sample in the database serves as the output (e.g., the federated median computation task). It realizes $\epsilon$-CDP by calculating a utility score (preference of serving as the output) for each data sample, and then selecting the output data sample based on the probabilistic distribution characterized by the utility scores. The exponential mechanism has been applied to realize CDP and also distributed differential privacy, which will be introduced later. [89] Conducting sampling on the data samples in the database is another effective tool to realize CDP. When only a random portion of data samples are selected for the data analytics task, randomness is introduced to the final analytics results, which can satisfy the criterion of CDP in (5) after proper mathematical proof. Two approaches are considered to utilize the sampling technique. In [18, 43], sampling is applied to directly realize CDP without any noise injection procedure; in [49, 55], sampling can reinforce a scheme that already satisfies DP via noise injection, by deriving the stronger privacy preservation (decreasing the $\epsilon$ parameter). The latter approach can be applied for both reinforcing CDP and LDP.
@@
Distributed differential privacy As mentioned above, LDP can achieve a stronger local privacy guarantee, without the assumption of a trusted aggregator; CDP introduces randomness with a smaller magnitude, which derives a higher correctness in final analytics results. Distributed differential privacy (DDP) is an attempt the combine the advantages of LDP and CDP. DDP statistically satisfies CDP to privatize the aggregated result. In addition, DDP requires clients to utilize cryptography tools to encrypt their uploads. With the encryption, the aforementioned local privacy risk existing in CDP, that an adversary can infer sensitive information when observing the individual client upload, is then eliminated.
@@
The deployment of DDP consists of two phases: distributed noise generation and encrypted aggregation. In the distributed noise generation phase, each client injects noise on their local uploads. The noise is deliberately designed so that the aggregated noise of the clients can satisfy the criterion of CDP. In the encrypted aggregation phase, the client upload, possessed by distributed noise generation and uploaded to the server with encryption. The server cannot observe the individual upload of any client, but can only observe the CDP-enforced aggregated uploads.
@@
$k$-anonymity The concept of $k$-anonymity emerged as a privacy-preserving technique, specifically crafted to safeguard individual data in publicly accessible data sets, addressing the growing concerns over personal data security and privacy breaches [91]. The main idea is to ensure that any specific individual cannot be linked given any record in the dataset by guaranteeing that any combination of attribute values will indicate at least $k$ individuals. In other words, for every individual in the data set, there are at least $k-1$ other members with the same information thus it’s rarely possible to search out an individual based solely on certain attribute values. To achieve $k$-anonymous, some attributes are either removed or suppressed. There are three types of attributes in datasets: identifiers, non-identifiers, and quasi-identifiers. Quasi-identifiers are attributes that can potentially distinguish individuals together such as age, zip code, and gender, none of which can identify an entity alone. Generally, identifiers are suppressed and quasi-identifier are either removed or generalized such as replacing the ages with age ranges to achieve $k$-anonymous.
@@
It’s a suitable tool in federated analytics in terms of preventing the leakage of sensitive information and the identification of individuals. Usually, $k$-anonymity is utilized in phases before sharing information. The risk of privacy breaches is reduced by guaranteeing that the data set of each client follows k-anonymity before sharing with the server or each other. In [46], they designed the Private Data Network to enable querying in databases of different parties without disclosure of their raw data by appling $k$-anonymity before data sharing. With the trusted third party as the query planner and results collector, the $k$-anonymous data of clients are then queried and aggregated to solve the questions from users. This setting is particularly relevant to collaboration in clinical research of healthcare clients. There are other methods to achieve $k$-anonymous. Researchers in [61] utilized sampling and trie pruning to fulfill $k$-anonymity for obtaining DP without noise injection. In [20], authors applied a more straightforward method that only the data sent by more than $k$ clients are chosen in the aggregation phase.
@@
Homomorphic encryption Homomorphic encryption is a powerful privacy-preserving cryptographic method allowing to perform mathematical computations such as addition and multiplication on encrypted data without decryption. It converts original data to cyphertexts by a public key, which cannot be read before it’s transformed back with the corresponding private key. Homomorphic encryption plays an important role in federated analytics in that it enables the aggregation of encrypted data without access to clients’ plaintext data. The collected encrypted data are then performed with mathematical operations to obtain a result which is also encrypted with support of its homomorphic attribute. The outcome is usually sent to a trusted third party holding the private key for decryption and is finally revealed to the querier. In this way, clients cooperate on data analysis tasks but raw data never leaves the local position of their owners. There are several types of homomorphic encryption, including fully and partially homomorphic encryption.
@@
Homomorphic encryption was early applied in private set intersection tasks[92, 93, 94, 95, 51]. Generally, in the two-party setting, the sender and receiver first agree on a homomorphic encryption scheme and the receiver generates a public-private key pair. The receiver then sends the encrypted set by public key to the sender who compares it with her encrypted set and sends the intersection back. Finally, the receiver revealed the encrypted intersection entities by the private key. The homomorphic attribute plays a crucial role in the faster computation of the phase in which the sender identifies common encrypted elements. Homomorphic encryption is also a suitable tool when there’s a requirement for allowed mathematical operations such as total summation in aggregation parts. In [29], researchers proposed Honeycrisp to solve the count mean sketch task at a large scale without a trusted core. A committee randomly selected from users generates keys and each client sends their additively homomorphically encrypted data by “two-element” Ring-LWE-based encryption to the server which obtains the total encrypted sum and sends it to the committee for decryption. Authors in [56] proposed FedMF, a secure matrix factorization framework that also used an additive homomorphic encryption, Paillier encryption, to secure the gradient information of each user.
TABLE IV: Summary of homomorphic encryption.
Reference Task Method Type
Freedman et al. [94] PSI Paillier encryption partially
Huang et al. [95] PSI ElGamal encryption partially
Freedman et al. [93] PSI ElGamal encryption partially
Chen et al. [92] PSI FHE with fine tuned parameters fully
Zhang et al. [51] PSI ElGamal encryption partially
Honeycrisp [29] Count mean sketch “two-element” Ring-LWE-based encryption fully
FedMF [56] Matrix factorization Paillier encryption partially
@@
Secure multi-party computation Secure multiparty computation (MPC) aims to jointly compute functions with inputs from several parties while guaranteeing the privacy of these in- puts. MPC ensures privacy and security in federated analytics by allowing multiple parties to collaborate on tasks without compromising the confidentiality of individual data sets of each party. Secure multi-party computation focuses less on the requirement for a large volume of data to maintain privacy and accuracy when compared to differential privacy. Nevertheless, it imposes a heavier burden in terms of computational and communication costs. While it can excel in preserving privacy and ensuring precise results with any data size, the trade-off lies in the heightened demand for computational resources and communication overhead among participated parties, rendering it a more resource-intensive approach. Some cryptographic techniques, like homomorphic encryption, garbled circuits, or secret sharing, are usually utilized to enable secure computation among parties. The choice of technique depends on the specific requirements and constraints of various federated analytics tasks. In the computation phase, only intermediate results are shared instead of users’ raw data, which are also usually encrypted to guarantee the throughout the whole process.
@@
MPC has a broad application in federated analytics. It’s used in tasks like secure communication of the median [30, 31], multi-party set intersection [96], and searching heavy hitters [44]. In [30], it focuses on securely computing the kth-ranked element (such as the median) of combined confidential datasets from multiple parties. The protocols employ techniques like binary search and consistency checks to ensure privacy while computing the median. While [31] addresses distributed private learning and focuses on computing the median in a way that is both differentially private and efficient in terms of computation and scalability. It utilizes a multi-party computation (MPC) approach to compute the exponential mechanism for the median, which is also adaptable for other rank-based statistics and machine learning optimizations. In [96], the authors introduce an efficient method for multi-party private set intersection (PSI), allowing multiple parties to compute the intersection of their datasets without revealing extra information. It uses a new paradigm based on oblivious programmable pseudorandom functions (OPPRF) and avoids computationally expensive public-key operations. Also, the protocol is secure against colluding semi-honest parties and has been demonstrated to be practical for up to 15 parties, each with datasets of a million items. Researchers in [44] present efficient multi-party computation (MPC) protocols, HH and PEM, to compute differentially private heavy hitters. These protocols use sketches for approximate counts, offering better accuracy than local differential privacy, and are practical for different data sizes: HH is suitable for small datasets (running time linear in data size), while PEM is more efficient for larger datasets (running time sublinear in data domain). The protocols are more accurate than local differential privacy approaches and an accessible performance on computation and communication cost. Also, some MPC protocols or platforms are designed to perform federated analytics [47, 97, 50, 51, 52]. Conclave [47] is a query compiler designed to accelerate relational analytics queries by combining data-parallel local cleartext processing with smaller secure multi-party computation (MPC) steps. It offers a hybrid MPC-cleartext protocol for cases where parties trust others with specific subsets of data. Cerebro [97] is an end-to-end collaborative learning platform that enables multiple parties to compute learning tasks without sharing plaintext data. It addresses the need of organizations to collaboratively use sensitive data, complying with policy regulations and business competition. Cerebro balances system design for safe collaboration with release policies and auditing, aiming to simplify the complex performance trade-offs between different MPC protocols. Authors in [50] introduce Senate, a system for MPC that allows multiple parties to collaboratively execute analytical SQL queries without exposing individual data, even with the presence of malicious adversaries. The main idea of Senate is MPC decomposition which enhances computational efficiency by breaking down cryptographic computations into smaller, parallel units. In [51], researchers focus on advancing the field of skyline queries, which are crucial for multi-criteria decision-making systems. It proposes a local dominance-based framework to enhance the efficiency of skyline queries in a vertical data federation setup. The framework decomposes skyline queries into more manageable units, improving the overall query process’s efficiency and security. Hu-Fu [52] is the first system dedicated to efficient and secure spatial query processing in a data federation context by optimizing the balance between plaintext and secure operations, minimizing the use of secure operators while still maintaining the overall security of the process. It’s a significant advancement in data federation that Hu-Fu parses federated spatial queries written in SQL, decomposes them into secure and plaintext components, and securely collects query results, offering a practical solution to the challenges of processing spatial queries securely and efficiently. Besides, authors in [29] designed a protocol to distribute the secret key of homomorphic encryption to members of a committee randomly composed of several clients. After the aggregator obtaining the total encrypted summation by public key from clients, it sent the result to the committee who then applied MPC to “fix” the secret the key and decrypted the result before revealing it. Functional encryption is another technique related to MPC. It expands upon public-key encryption by allowing an individual with a secret key to decipher a particular function derived from the encrypted content in the ciphertext. In [98], authors proposed the FAA-DL where clients and the server can collaboratively and proactively analyze anomaly based on functional encryption.

### v3

@@ V-A2 Central differential privacy
Compared to LDP, where the client data are considered data samples and directly protected, CDP cannot be directly applied to many FA scenarios because it considers enforcing protection on the data already aggregated. In the FA literature, there are mainly two settings to enforce CDP. In the first setting, the raw data of each client is considered as a database consisting of multiple data samples. The clients enforce CDP on the database, so that information about each sample is protected. FA research considering this setting includes [44, 77]. In the second setting, similar to the setting of LDP, the data within a single client is considered as a data sample in $\mathcal{D}$. The CDP scheme guarantees that an adversary cannot infer the sensitive information of a client by analyzing the final FA result. However, such privacy attacks can succeed when the adversary has access to the uploads of the clients (e.g., when the server acts as the adversary). Therefore, the second setting usually assumes a trusted aggregator (server) and sacrifices the formal privacy guarantee. The second setting has been considered in [42, 43, 28, 70].
@@
These exist diverse techniques to realize CDP. Similar to those used for LDP, the Gaussian mechanism can still be applied for CDP to protect continuous scaler output, which can satisfy $(\epsilon,\delta)$-CDP [44, 77, 42, 43]. [^4] Laplace mechanism also has theoretically applicable for CDP, but is not selected for researchers due to the high variance of Laplace noise. Compared to the LDP case, where Gaussian noise is added to each single data sample, applying the Gaussian mechanism for CDP only requires adding noise to the aggregated result. Therefore, the overall noise on the final analytics results is significantly reduced. The exponential mechanism tackles the analytics tasks where a single sample in the database serves as the output (e.g., the federated median computation task). It realizes $\epsilon$-CDP by calculating a utility score (preference of serving as the output) for each data sample, and then selecting the output data sample based on the probabilistic distribution characterized by the utility scores. The exponential mechanism has been applied to realize CDP and also DDP, which will be introduced later. [51] Conducting sampling on the data samples in the database is another effective tool to realize CDP. When only a random portion of data samples are selected for the data analytics task, randomness is introduced to the final analytics results, which can satisfy the criterion of CDP in (5) after proper mathematical proof. Two approaches are considered to utilize the sampling technique. In [28, 70], sampling is applied to directly realize CDP without any noise injection procedure; in [77, 83], sampling can reinforce a scheme that already satisfies DP via noise injection, by deriving the stronger privacy preservation (decreasing the $\epsilon$ parameter). The latter approach can be applied for both reinforcing CDP and LDP. This approach to realizing CDP is novel as an FA mechanism because it can be realized only when the sampling of clients is available. It is not appropriate for conventional centralized and distributed algorithms where the sampling of clients is not available.
Lessons learned. CDP assumes that the raw data are already gathered by an aggregator, and the noise helps protect individual data records when the aggregated statistics are released. Since conventional FA settings usually do not assume a trusted aggregator, and the clients should preserve data privacy against the server, FA solutions utilizing CDP essentially sacrifice privacy preservation to some extent. Compared to LDP, CDP requires less noise to the uploads and even can be realized via sampling without adding noise to the raw value. CDP can gain higher data utility under weaker adversarial settings.
V-A3 Distributed differential privacy
@@
As mentioned above, LDP can achieve a stronger local privacy guarantee, without the assumption of a trusted aggregator; CDP introduces randomness with a smaller magnitude, which derives a higher correctness in final analytics results. DDP is an attempt the combine the advantages of LDP and CDP, or an attempt to apply CDP in the local privacy model. DDP statistically satisfies CDP to privatize the aggregated result. In addition, DDP requires clients to utilize cryptography tools to encrypt their uploads. With the encryption, the aforementioned local privacy risk existing in CDP, that an adversary can infer sensitive information when observing the individual client upload, is then eliminated. The idea of DDP is inspired by [137], and the primary formal DDP designs is proposed in [138, 139], which are much earlier than the introduction of the term “DDP”. Essentially, DDP describes the idea of extending the privacy preservation of CDP to the local model in the distributed setting with the help of cryptography tools.
@@
The deployment of DDP consists of two phases: distributed noise generation and encrypted aggregation. In the distributed noise generation phase, each client injects noise into their local uploads. The noise is deliberately designed so that the aggregated noise of the clients can satisfy the criterion of CDP. In the encrypted aggregation phase, the client upload is possessed by distributed noise generation and uploaded to the server with encryption. The server cannot observe the individual upload of any client, but can only observe the CDP-enforced aggregated uploads.
@@
Lessons learned. DDP utilizes cryptography tools to force an aggregator unable to observe raw uploads during aggregation so that the magnitude of required noise can be reduced, which provides an option of privacy preservation with higher data utility (equivalent to CDP) in the typical FA setting without a trusted aggregator. However, CDP introduces cryptography tools into its design realm, making DDP mechanisms have much higher computation and communication overheads, compared to conventional DP mechanisms.
@@
DP has proven its effectiveness in privacy preservation from the statistical perspective. However, it is sometimes hard to design mechanisms to realize the DP criteria defined in (4) or (5), especially in cases when the privatized data is in high dimensionality or complex structure. Therefore, researchers design relaxations of DP that are easier to satisfy and meet particular privacy needs. Examples of these DP criteria include Rényi DP [141], concentrated DP [142], and dependent DP [143]. Variations of DP are also introduced to handle specific data types with specific definitions of adjacent datasets or units of privacy preservation, such as sample label [144], streaming data [145], graph data [146], location data [147], and natural language data [148]. Generalizations of DP, like pufferfish privacy [149] and blowfish privacy [150], are also introduced to handle the limitations of DP.
@@ V-A4 $k$-anonymity
The concept of $k$-anonymity emerged as a privacy-preserving technique, specifically crafted to safeguard individual data in publicly accessible datasets, addressing the growing concerns over personal data security and privacy breaches [151, 152, 153]. The main idea is to ensure that any specific individual cannot be linked given any record in the dataset by guaranteeing that any combination of attribute values will indicate at least $k$ individuals. In other words, for every individual in the dataset, there are at least $k-1$ other members with the same information thus it is rarely possible to search out an individual based solely on certain attribute values. To achieve $k$-anonymous, some attributes are either removed or suppressed. There are three types of attributes in datasets: identifiers, non-identifiers, and quasi-identifiers. Quasi-identifiers are attributes that can potentially distinguish individuals together such as age, zip code, and gender, none of which can identify an entity alone. Generally, identifiers are suppressed and quasi-identifiers are either removed or generalized such as replacing the ages with age ranges to achieve $k$-anonymous.
@@
It is a suitable tool in FA in terms of preventing the leakage of sensitive information and the identification of individuals. Usually, $k$-anonymity is utilized in phases before sharing information. The risk of privacy breaches is reduced by guaranteeing that the dataset of each client follows $k$-anonymity before sharing with the server or each other. In [74], they designed the Private Data Network to enable querying in databases of different parties without disclosure of their raw data by applying $k$-anonymity before data sharing. With the trusted third party as the query planner and results collector, the $k$-anonymous data of clients are then queried and aggregated to solve the questions from users. This setting is particularly relevant to collaboration in clinical research of healthcare clients. There are other methods to achieve $k$-anonymous. Researchers in [125] utilized sampling and tree pruning to fulfill $k$-anonymity for obtaining DP without noise injection. In [30], authors applied a more straightforward method that only the data sent by more than $k$ clients are chosen in the aggregation phase.
Lessons learned. While k-anonymity is a versatile privacy-preserving technique, its application in FA is uniquely suited to address the challenges of data aggregation and querying without compromising individual privacy. The specific methods and practical implementations of k-anonymity are uniquely applied in FA, such as pre-sharing anonymization and the use of trusted third parties. However, k-anonymity has notable limitations that can affect its efficacy. It is vulnerable to attribute linkage attacks, where external knowledge can be used to re-identify individuals, and to homogeneity attacks, where all records in a group have the same sensitive attribute, making them easily identifiable. Achieving k-anonymity often leads to significant data utility loss due to generalization and suppression techniques, which is further exacerbated in datasets with many quasi-identifiers because of the curse of dimensionality.
V-A5 Homomorphic encryption
@@
Homomorphic encryption (HE) is a powerful privacy-preserving cryptographic method allowing clients to perform mathematical computations such as addition and multiplication on encrypted data [154]. The output of HE computation is still in the encrypted form, and when the output is decrypted, the result will be identical to the corresponding computation results on encrypted raw data. HE plays an important role in FA in that it enables the aggregation of encrypted data without access to clients’ plaintext data. The collected encrypted data are then performed with mathematical operations to obtain a result which is also encrypted with support of its homomorphic attribute. The outcome is usually sent to a trusted third party holding the private key for decryption and is finally revealed to the querier. In this way, clients cooperate on data analysis tasks but raw data never leaves the local position of their owners. There are several types of HE, including fully [155, 156, 157] and partially [158, 159, 160] HE. Specific implementations in FA, such as private set intersection and large-scale data aggregation without a trusted core, showcase its crucial role in privacy-preserving analytics. In contrast, in FL, HE is used more for securing model updates during training rather than for data aggregation, highlighting its distinct and central role in FA.
@@
HE was early applied in private set intersection tasks[161, 162, 163, 164, 79]. Generally, in the two-party setting, the sender and receiver first agree on an HE scheme and the receiver generates a public-private key pair. The receiver then sends the encrypted set by public key to the sender who compares it with her encrypted set and sends the intersection back. Finally, the receiver revealed the encrypted intersection entities by the private key. The homomorphic attribute plays a crucial role in the faster computation of the phase in which the sender identifies common encrypted elements. HE is also a suitable tool when there’s a requirement for allowed mathematical operations such as total summation in aggregation parts. In [37], researchers proposed Honeycrisp to solve the count mean sketch task at a large scale without a trusted core. A committee randomly selected from users generates keys and each client sends their additively homomorphically encrypted data by “two-element” Ring-LWE-based encryption to the server which obtains the total encrypted sum and sends it to the committee for decryption.
Lessons learned. HE is a powerful cryptographic technique that enables mathematical operations on encrypted data without decryption, ensuring privacy in data analysis. In FA, HE allows secure aggregation of encrypted data from multiple clients without exposing their raw data, with the final result sent to a trusted third party for decryption. HE plays a crucial role in privacy-preserving tasks such as private set intersection and large-scale data aggregation, offering strong security guarantees. While its use in FL focuses on securing model updates, in FA, HE primarily supports secure data aggregation, highlighting its distinct roles in different privacy-preserving contexts. However, HE has notable drawbacks that can hinder its practical adoption. Operations on encrypted data are computationally intensive, leading to high processing times, especially for complex computations, and the encrypted data often requires significantly more storage and communication bandwidth due to large ciphertext sizes. Additionally, while Fully Homomorphic Encryption supports arbitrary computations, it is resource-intensive and less practical for real-time applications, unlike Partial Homomorphic Encryption, which has limited functionality. Furthermore, managing and securely distributing encryption keys adds complexity, making implementation and maintenance challenging.
V-A6 Secure multi-party computation
@@
Secure multi-party computation (MPC) aims to jointly compute functions with inputs from several parties while guaranteeing the privacy of these inputs [165]. MPC allows the computation with input held by multiple parties, where the result of the computation result can be derived while the input data are held private. MPC ensures privacy and security in FA by allowing multiple parties to collaborate on tasks without compromising the confidentiality of individual data sets of each party, utilizing the technologies of oblivious transfer and secret sharing. MPC focuses less on the requirement for a large volume of data to maintain privacy and accuracy when compared to DP. Nevertheless, it imposes a heavier burden in terms of computational and communication costs [166, 167]. While it can excel in preserving privacy and ensuring precise results with any data size, the trade-off lies in the heightened demand for computational resources and communication overhead among participated parties, rendering it a more resource-intensive approach. Some cryptographic techniques, like HE, garbled circuits, or secret sharing, are usually utilized to enable secure computation among parties. The choice of technique depends on the specific requirements and constraints of various FA tasks. In the computation phase, only intermediate results are shared instead of users’ raw data, which are also usually encrypted to guarantee the throughout the whole process.
@@
Some researchers directly use MPC as a privatization tool to encrypt the computation results and provide privacy preservation for certain analytics algorithms. MPC is used in tasks like secure communication of the median [40, 41], multi-party set intersection [168], and searching heavy hitters [71]. In [40], it focuses on securely computing the k-th-ranked element (such as the median) of combined confidential datasets from multiple parties. The protocols employ techniques like binary search and consistency checks to ensure privacy while computing the median. While [41] addresses distributed private learning and focuses on computing the median in a way that is both differentially private and efficient in terms of computation and scalability. It utilizes an MPC approach to compute the exponential mechanism for the median, which is also adaptable for other rank-based statistics and machine learning optimizations. In [168], the authors introduce an efficient method for multi-party private set intersection (PSI), allowing multiple parties to compute the intersection of their datasets without revealing extra information. It uses a new paradigm based on oblivious programmable pseudorandom functions (OPPRF) and avoids computationally expensive public-key operations. Also, the protocol is secure against colluding semi-honest parties and has been demonstrated to be practical for up to 15 parties, each with datasets of a million items. Researchers in [71] present efficient MPC protocols, HH and PEM, to compute differentially private heavy hitters. These protocols use sketches for approximate counts, offering better accuracy than LDP, and are practical for different data sizes: HH is suitable for small datasets (running time linear in data size), while PEM is more efficient for larger datasets (running time sublinear in data domain). The protocols are more accurate than LDP approaches and have an accessible performance on computation and communication costs.
In addition, some researchers design MPC protocols and platforms for FA applications, rather than directly using MPC to complete certain FA tasks [75, 169, 78, 79, 80]. The protocol & platform-based approach provides the users the capacity to apply the MPC scheme to customized tasks and raises the possibility of multipurpose FA systems. Conclave [75] is a query compiler designed to accelerate relational analytics queries by combining data-parallel local cleartext processing with smaller MPC steps. It offers a hybrid MPC-cleartext protocol for cases where parties trust others with specific subsets of data. Cerebro [169] is an end-to-end collaborative learning platform that enables multiple parties to compute learning tasks without sharing plaintext data. It addresses the need of organizations to collaboratively use sensitive data, comply with policy regulations, and deal with business competition. Cerebro balances system design for safe collaboration with release policies and auditing, aiming to simplify the complex performance trade-offs between different MPC protocols. Authors in [78] introduce Senate, a system for MPC that allows multiple parties to collaboratively execute analytical SQL queries without exposing individual data, even with malicious adversaries. The main idea of Senate is MPC decomposition which enhances computational efficiency by breaking down cryptographic computations into smaller, parallel units. In [79], researchers focus on advancing the field of skyline queries, which are crucial for multi-criteria decision-making systems. It proposes a local dominance-based framework to enhance the efficiency of skyline queries in a vertical data federation setup. The framework decomposes skyline queries into more manageable units, improving the overall query process’s efficiency and security. Hu-Fu [80] is the first system dedicated to efficient and secure spatial query processing in a data federation context by optimizing the balance between plaintext and secure operations, minimizing the use of secure operators while maintaining the process’s overall security. It’s a significant advancement in data federation that Hu-Fu parses federated spatial queries written in SQL, decomposes them into secure and plaintext components, and securely collects query results, offering a practical solution to the challenges of processing spatial queries securely and efficiently. Besides, authors in [37] designed a protocol to distribute the secret key of HE to members of a committee randomly composed of several clients. After the aggregator obtains the total encrypted summation by public key from clients, it sends the result to the committee who then applies MPC to “fix” the secret key and decrypt the result before revealing it. Functional encryption is another technique related to MPC. It expands upon public-key encryption by allowing an individual with a secret key to decipher a particular function derived from the encrypted content in the ciphertext. In [170], authors proposed the FAA-DL where clients and the server can collaboratively and proactively analyze anomalies based on functional encryption.
Lessons learned. MPC plays a pivotal role in ensuring privacy and security in FA by enabling multiple parties to collaborate without revealing their datasets. MPC ensures that only intermediate results are shared, and the raw data remains private throughout the process. However, it comes with a significant trade-off in terms of computational and communication costs, which can be resource-intensive. Despite this, MPC excels in preserving privacy and delivering accurate results, regardless of data size, making it suitable for tasks such as secure set intersection, median computation, and heavy hitters search. Furthermore, the development of MPC protocols and platforms has paved the way for more efficient, flexible, and scalable FA systems, allowing users to apply MPC to various customized tasks. These systems balance the complexities of privacy preservation with system performance, offering practical solutions for secure collaborative data analysis across different domains.